# Tutorial on amortized optimization

### v4

@@ 4.1.3 Control with a model-based value estimate
Extending the results of section 4.1.2, this section compares the trained humanoid policy from https://github.com/facebookresearch/svg to solving a short-horizon ($H=3$) model-based control optimization problem defined in eq. 3.38. The optimal action sequence solving eq. 3.38 is $u^{\star}_{1:H}$ can be approximated by interleaving a model-free policy $\pi_{\theta}$ with the dynamics $f$. While standard model predictive control method are often ideal for solving for $u^{\star}_{1:H}$ from scratch, using Adam as a gradient-based shooting method is a reasonable baseline in this short-horizon setting.
Figure 4.6 shows that the amortized policy is approximately 700 times faster than solving the problem from scratch. This model-based setting has the same issues with the approximation errors in the models and the model-based value estimate is again normalized and clamped so that the expert policy has a value of zero and the zero action has a value of -1. The source code behind this example is also available in code/evaluate_amortization_speed_function_control.py.
@@
Figure 4.6: Runtime comparison between Adam and a learned policy $\pi_{\theta}$ to solve a short-horizon ($H=3$) model-based control problem (eq. 3.38) on the humanoid MDP. This was evaluated as a batch on an expert trajectory with 1000 states and was run on an unloaded NVIDIA Quadro GP100 GPU. The amortized policy is approximately 700 times faster than solving the problem from scratch.
@@ 4.2 Training an amortization model on a sphere
where ${\mathcal{S}}^{2}$ is the surface of the unit 2-sphere embedded in $\mathbb{R}^{3}$ as ${\mathcal{S}}^{2}\vcentcolon=\{y\in\mathbb{R}^{3}\mid\|y\|_{2}=1\}$ and $x$ is some parameterization of the function $f:{\mathcal{S}}^{2}\times{\mathcal{X}}\rightarrow\mathbb{R}$. Eq. 4.1 is relevant to physical and geographical settings seeking the extreme values of a function defined on the Earth or other spaces that can be approximated with a sphere. The full source code behind this experiment is available in code/train-sphere.py.
Amortization objective. Eq. 4.1 first needs to be transformed from a constrained optimization problem into an unconstrained one of the form eq. 1.1. In this setting, one way of doing this is by using a projection:
@@
$\pi_{{\mathcal{S}}^{2}}(x)\vcentcolon=$ $\operatorname*{arg\,min}_{y\in{\mathcal{S}}^{2}}\|y-x\|_{2}$ (4.3)
@@
Amortization model. The model $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow\mathbb{R}$ is a fully-connected MLP. The predictions to eq. 4.1 on the sphere can again be obtained by composing the output with the projection $\pi_{{\mathcal{S}}^{2}}\circ\hat{y}_{\theta}$.
Optimizing the gradient-based loss. Finally, it is reasonable to optimize the gradient-based loss ${\mathcal{L}}_{\rm obj}$ because the objective and model are tractable and easily differentiable. Figure 4.7 shows the model’s predictions starting with the untrained model and finishing with the trained model, showing that this setup indeed enables us to predict the solutions to eq. 4.1 with a single neural network $\hat{y}_{\theta}(x)$ trained with the gradient-based loss.
@@ 4.3 Other useful software packages
  • higher [Grefenstette et al., 2019] is a PyTorch library that adds differentiable higher-order optimization support with 1) monkey-patched functional torch.nn modules, and 2) differentiable versions of torch.optim optimizers such as Adam and SGD. This enables arbitrary torch modules and optimizers to be unrolled and used as as a semi-amortized model.
@@ 5.1 Surpassing the convergence rates of classical methods
Theoretical and empirical optimization research often focuses on discovering algorithms with theoretically strong convergence rates in general or worst-case scenarios. Many of the algorithms with the best convergence rates are used as the state-of-the-art algorithms in practice, such as momentum and acceleration methods. Amortized optimization methods can surpass the results provided by classical optimization methods because they are capable of tuning the initialization and updates to the best-case scenario within the distribution of contexts the amortization model is trained on. For example, the fully amortized models for amortized variational inference and model-free actor-critic methods for RL presented in section 4.1 solve the optimization problems in constant time with just a single prediction of the solution from the context without even looking at the objective! Further theoretical characterizations of this are provided in Khodak et al. [2022] and related literature on algorithms with predictions.
@@ 5.3 Measuring performance
Quantifying the performance of amortization models can be even more challenging than the choice between using a regression- or objective-based loss and is often tied to problem-specific metrics that are important. For example, even if a method is able to attain low objective values in a few iterations, the computation may take longer than a specialized algorithm or another amortization model that can reach the same level of accuracy, thus not making it useful for the original goal of speeding up solves to eq. 1.1.
@@ 5.4 Successes and limitations of amortized optimization
  • Objective $f(y;x)$ is smooth over the domain ${\mathcal{Y}}$ and has unique solutions $y^{\star}$. With objective-based learning, non-convex objectives with few poor local optima are ideal. This behavior can be encouraged with smoothing as is often done for meta-learning and policy learning (section 2.2.3).
  • A higher-level process should tolerate sub-optimal solutions given by $\hat{y}$ in the beginning of training. In variational encoders, the suboptimal bound on the likelihood is still acceptable to optimize the density model’s parameters over in eq. 3.5. And in reinforcement learning policies, a suboptimal solution to the maximum value problem is still acceptable to deploy on the system in early phases of training, and may even be desirable for the exploration induced by randomly initialized policies.
@@
  • The domain requires accurate solutions. Even though metrics that measure the solution quality of $\hat{y}$ can be defined on top of eq. 1.1, amortized methods typically cannot rival the accuracy of standard algorithms used to solve the optimization problems. In these settings, amortized optimization still has the potential at uncovering new foundations and algorithms for solving problems, but is non-trivial to successfully demonstrate. From an amortization perspective, one difficulty of safety-critical model-free reinforcement learning comes from needing to ensure the amortized policy properly optimizes a value estimate that (hopefully) encodes safety-critical properties of the state-action space.
@@ 5.5 Some open problems and under-explored directions
    Overcoming local minima with objective-based losses and connections to stochastic policies. Section 2.2.3 covered the objective smoothing by Metz et al. [2019a], Merchant et al. [2021] to overcome suboptimal local minima in the objective. These have striking similarities to stochastic policies in reinforcement learning that also overcome local minima, e.g. in eq. 3.39. The stochastic policies, such as in Haarnoja et al. [2018], have the desirable property of starting with a high variance and then focusing in on a low-variance solution with a penalty constraining the entropy to a fixed value. A similar method is employed in GECO [Rezende and Viola, 2018] that adjusts a Lagrange multiplier in the ELBO objective to achieve a target conditional log-likelihood. These tricks seem useful to generalize and apply to other amortization settings to overcome poor minima.
@@
    Distribution shift of $p(x)$ and out-of-distribution generalization. This tutorial has assumed that $p(x)$ is fixed and remains the same through the entire training process. However, in some settings $p(x)$ may shift over time, which could come from 1) the data generating process naturally changing, or 2) a higher-level learning process also influencing $p(x)$. Furthermore, after training on some context distribution $p(x)$, a deploy model is likely not going to be evaluated on the same distribution and should ideally be resilient to out-of-distribution samples. The out-of-distribution performance can often be measured and quantified and reported alongside the model. Even if the amortization model fails at optimizing eq. 1.1, it’s detectable because the optimality conditions of eq. 1.1 or other solution quality metrics can be checked. If the solution quality isn’t high enough, then a slower optimizer could potentially be used as a fallback.
@@
    Amortized and semi-amortized control and reinforcement learning. Applications of semi-amortization in control and reinforcement learning covered in section 3.6 are budding and learning sample-efficient optimal controllers is an active research area, especially in model-based settings where the dynamics model is known or approximated. Amos and Yarats [2020] shows how amortization can learn latent control spaces that are aware of the structure of the solutions to control problems. Marino et al. [2021] study semi-amortized methods based on gradient descent and show that they better-amortize the solutions than the standard fully-amortized models.
@@ 5.6.2 Amortized optimization over discrete domains
A significant generalization of eq. 1.1 is to optimization problems that have discrete domains, which includes combinatorial optimization and mixed discrete-continuous optimization. I have chosen to not include these works in this tutorial as many methods for discrete optimization are significantly different from the methods considered here, as learning with derivative information often becomes impossible. Key works in discrete and combinatorial spaces include Khalil et al. [2016, 2017], Jeong and Song [2019], Bertsimas and Stellato [2019], Shao et al. [2021], Bertsimas and Stellato [2021], Cappart et al. [2021] and the surveys [Lodi and Zarpellon, 2017, Bengio et al., 2021, Kotary et al., 2021] capture a much broader view of this space. Banerjee and Roy [2015] consider repeated ILP solves and show applications in aircraft carrier deck scheduling and vehicle routing. For architecture search, Luo et al. [2018] learn a continuous latent space behind the discrete architecture space. Many reinforcement learning and control methods over discrete spaces can also be seen as amortizing or semi-amortizing the discrete control problems, for example: Cauligi et al. [2020, 2021] use regression-based amortization to learn mixed-integer control policies. Fickinger et al. [2021] fine-tune the policy optimizer for every encountered state. Tennenholtz and Mannor [2019], Chandak et al. [2019], Van de Wiele et al. [2020] learn latent action spaces for high-dimensional discrete action spaces with shared structure.

### v5

@@ 4.1.3 Control with a model-based value estimate
Extending the results of section˜4.1.2, this section compares the trained humanoid policy from https://github.com/facebookresearch/svg to solving a short-horizon ($H=3$) model-based control optimization problem defined in eq.˜3.38. The optimal action sequence solving eq.˜3.38 is $u^{\star}_{1:H}$ can be approximated by interleaving a model-free policy $\pi_{\theta}$ with the dynamics $f$. While standard model predictive control method are often ideal for solving for $u^{\star}_{1:H}$ from scratch, using Adam as a gradient-based shooting method is a reasonable baseline in this short-horizon setting.
Figure˜4.6 shows that the amortized policy is approximately 700 times faster than solving the problem from scratch. This model-based setting has the same issues with the approximation errors in the models and the model-based value estimate is again normalized and clamped so that the expert policy has a value of zero and the zero action has a value of -1. The source code behind this example is also available in code/evaluate_amortization_speed_function_control.py.
@@
Figure 4.6: Runtime comparison between Adam and a learned policy $\pi_{\theta}$ to solve a short-horizon ($H=3$) model-based control problem (eq.˜3.38) on the humanoid MDP. This was evaluated as a batch on an expert trajectory with 1000 states and was run on an unloaded NVIDIA Quadro GP100 GPU. The amortized policy is approximately 700 times faster than solving the problem from scratch.
@@ 4.2 Training an amortization model on a sphere
where ${\mathcal{S}}^{2}$ is the surface of the unit 2-sphere embedded in $\mathbb{R}^{3}$ as ${\mathcal{S}}^{2}\vcentcolon=\{y\in\mathbb{R}^{3}\mid\|y\|_{2}=1\}$ and $x$ is some parameterization of the function $f:{\mathcal{S}}^{2}\times{\mathcal{X}}\rightarrow\mathbb{R}$. Eq.˜4.1 is relevant to physical and geographical settings seeking the extreme values of a function defined on the Earth or other spaces that can be approximated with a sphere. The full source code behind this experiment is available in code/train-sphere.py.
Amortization objective. Eq.˜4.1 first needs to be transformed from a constrained optimization problem into an unconstrained one of the form eq.˜1.1. In this setting, one way of doing this is by using a projection:
@@
$\pi_{{\mathcal{S}}^{2}}(x)$= $\operatorname*{arg\,min}_{y\in{\mathcal{S}}^{2}}\|y-x\|_{2}$ (4.3)
@@
Amortization model. The model $\hat{y}_{\theta}:{\mathcal{X}}\rightarrow\mathbb{R}$ is a fully-connected MLP. The predictions to eq.˜4.1 on the sphere can again be obtained by composing the output with the projection $\pi_{{\mathcal{S}}^{2}}\circ\hat{y}_{\theta}$.
Optimizing the gradient-based loss. Finally, it is reasonable to optimize the gradient-based loss ${\mathcal{L}}_{\rm obj}$ because the objective and model are tractable and easily differentiable. Figure˜4.7 shows the model’s predictions starting with the untrained model and finishing with the trained model, showing that this setup indeed enables us to predict the solutions to eq.˜4.1 with a single neural network $\hat{y}_{\theta}(x)$ trained with the gradient-based loss.
@@ 4.3 Other useful software packages
  • higher [Grefenstette et al., 2019] is a PyTorch library that adds differentiable higher-order optimization support with 1) monkey-patched functional torch.nn modules, and 2) differentiable versions of torch.optim optimizers such as Adam and SGD. This enables arbitrary torch modules and optimizers to be unrolled and used as a semi-amortized model.
@@ 5.1 Surpassing the convergence rates of classical methods
Theoretical and empirical optimization research often focuses on discovering algorithms with theoretically strong convergence rates in general or worst-case scenarios. Many of the algorithms with the best convergence rates are used as the state-of-the-art algorithms in practice, such as momentum and acceleration methods. Amortized optimization methods can surpass the results provided by classical optimization methods because they are capable of tuning the initialization and updates to the best-case scenario within the distribution of contexts the amortization model is trained on. For example, the fully amortized models for amortized variational inference and model-free actor-critic methods for RL presented in section˜4.1 solve the optimization problems in constant time with just a single prediction of the solution from the context without even looking at the objective! Further theoretical characterizations of this are provided in Khodak et al. [2022] and related literature on algorithms with predictions.
@@ 5.3 Measuring performance
Quantifying the performance of amortization models can be even more challenging than the choice between using a regression- or objective-based loss and is often tied to problem-specific metrics that are important. For example, even if a method is able to attain low objective values in a few iterations, the computation may take longer than a specialized algorithm or another amortization model that can reach the same level of accuracy, thus not making it useful for the original goal of speeding up solves to eq.˜1.1.
@@ 5.4 Successes and limitations of amortized optimization
  • Objective $f(y;x)$ is smooth over the domain $\mathcal{Y}$} and has unique solutions $y^{\star}$. With objective-based learning, non-convex objectives with few poor local optima are ideal. This behavior can be encouraged with smoothing as is often done for meta-learning and policy learning (section˜2.2.3).
  • A higher-level process should tolerate sub-optimal solutions given by $\hat{y}$ in the beginning of training. In variational encoders, the suboptimal bound on the likelihood is still acceptable to optimize the density model’s parameters over in eq.˜3.5. And in reinforcement learning policies, a suboptimal solution to the maximum value problem is still acceptable to deploy on the system in early phases of training, and may even be desirable for the exploration induced by randomly initialized policies.
@@
  • The domain requires accurate solutions. Even though metrics that measure the solution quality of $\hat{y}$ can be defined on top of eq.˜1.1, amortized methods typically cannot rival the accuracy of standard algorithms used to solve the optimization problems. In these settings, amortized optimization still has the potential at uncovering new foundations and algorithms for solving problems, but is non-trivial to successfully demonstrate. From an amortization perspective, one difficulty of safety-critical model-free reinforcement learning comes from needing to ensure the amortized policy properly optimizes a value estimate that (hopefully) encodes safety-critical properties of the state-action space.
@@ 5.5 Some open problems and under-explored directions
    Overcoming local minima with objective-based losses and connections to stochastic policies. Section˜2.2.3 covered the objective smoothing by Metz et al. [2019a], Merchant et al. [2021] to overcome suboptimal local minima in the objective. These have striking similarities to stochastic policies in reinforcement learning that also overcome local minima, e.g. in eq.˜3.39. The stochastic policies, such as in Haarnoja et al. [2018], have the desirable property of starting with a high variance and then focusing in on a low-variance solution with a penalty constraining the entropy to a fixed value. A similar method is employed in GECO [Rezende and Viola, 2018] that adjusts a Lagrange multiplier in the ELBO objective to achieve a target conditional log-likelihood. These tricks seem useful to generalize and apply to other amortization settings to overcome poor minima.
@@
    Distribution shift of $p(x)$ and out-of-distribution generalization. This tutorial has assumed that $p(x)$ is fixed and remains the same through the entire training process. However, in some settings $p(x)$ may shift over time, which could come from 1) the data generating process naturally changing, or 2) a higher-level learning process also influencing $p(x)$. Furthermore, after training on some context distribution $p(x)$, a deploy model is likely not going to be evaluated on the same distribution and should ideally be resilient to out-of-distribution samples. The out-of-distribution performance can often be measured and quantified and reported alongside the model. Even if the amortization model fails at optimizing eq.˜1.1, it’s detectable because the optimality conditions of eq.˜1.1 or other solution quality metrics can be checked. If the solution quality isn’t high enough, then a slower optimizer could potentially be used as a fallback.
@@
    Amortized and semi-amortized control and reinforcement learning. Applications of semi-amortization in control and reinforcement learning covered in section˜3.6 are budding and learning sample-efficient optimal controllers is an active research area, especially in model-based settings where the dynamics model is known or approximated. Amos and Yarats [2020] shows how amortization can learn latent control spaces that are aware of the structure of the solutions to control problems. Marino et al. [2021] study semi-amortized methods based on gradient descent and show that they better-amortize the solutions than the standard fully-amortized models.
@@ 5.6.2 Amortized optimization over discrete domains
A significant generalization of eq.˜1.1 is to optimization problems that have discrete domains, which includes combinatorial optimization and mixed discrete-continuous optimization. I have chosen to not include these works in this tutorial as many methods for discrete optimization are significantly different from the methods considered here, as learning with derivative information often becomes impossible. Key works in discrete and combinatorial spaces include Khalil et al. [2016, 2017], Jeong and Song [2019], Bertsimas and Stellato [2019], Shao et al. [2021], Bertsimas and Stellato [2021], Cappart et al. [2021] and the surveys [Lodi and Zarpellon, 2017, Bengio et al., 2021, Kotary et al., 2021] capture a much broader view of this space. Banerjee and Roy [2015] consider repeated ILP solves and show applications in aircraft carrier deck scheduling and vehicle routing. For architecture search, Luo et al. [2018] learn a continuous latent space behind the discrete architecture space. Many reinforcement learning and control methods over discrete spaces can also be seen as amortizing or semi-amortizing the discrete control problems, for example: Cauligi et al. [2020, 2021] use regression-based amortization to learn mixed-integer control policies. Fickinger et al. [2021] fine-tune the policy optimizer for every encountered state. Tennenholtz and Mannor [2019], Chandak et al. [2019], Van de Wiele et al. [2020] learn latent action spaces for high-dimensional discrete action spaces with shared structure.